OpenStudy (anonymous):

Find the distance between the points given. (-3, 0) and (0, radical 7)

OpenStudy (anonymous):

can't remember how 2 do this XD

OpenStudy (anonymous):

Recall or remember, what the formula for finding the distance between two points is, that formula goes as follows\[d=\sqrt{(X_2-X_1)^2+(Y_2-Y_1)^2}\]now what does this mean? it means you take any two points, you label one of them to be\[(X_1,Y_1)\]and the other to be\[(X_2,Y_2)\]and just plug it into that formula, that's a very important formula, try to memorize it.

OpenStudy (anonymous):

I know that part, but the answer I got wasn't one of my choices.

OpenStudy (anonymous):

alrighty, gimme a second to do the calculations then, and then we can see if we got the same anser

OpenStudy (anonymous):

you should have gotten positive and negative 4, idk whether you need both answers or not, but that's what i got

OpenStudy (anonymous):

did u get 25 as an answer

OpenStudy (anonymous):

I got 9 plus sqrt 49

OpenStudy (anonymous):

must have done mines differently or wrong XD

OpenStudy (anonymous):

\[\sqrt{(-3-0)^2+(0-\sqrt{7})^2}=\sqrt{(-3)^2+(-\sqrt{7}^2)}\]so since you're squaring the whole thing in the parenthesis, they become positive\[\sqrt{9+7}=\sqrt{16}=\pm4\]

OpenStudy (anonymous):

Oh I see what i did wrong. Thank you. Could you help me with this one too? Find the distance between the points given: (-10,3) and (-10,12)

OpenStudy (anonymous):

same idea, pick one set of points to be your "set 1" and the other becomes "set 2" and plug them in the function, so you'd have to solve\[d=\sqrt{(-10-(-10))^2+(12-3)^2}=\sqrt{(0)^2+(9)^2}\]you can finish it, but just remember, to pick one set of coordinates to be the set 1, otherwise you could get the wrong answer, and another note, just remember that when you subtract a negative number, the negatives cancel and become a plus or positive

OpenStudy (anonymous):

Why wouldn't it be 3-12 sqrd for the second part

OpenStudy (anonymous):

For me, since i know the formula so well, i can notice that i don't really want to deal with\[(3-12)^2\]which would be negative, sure the result is the same, but really, it's because i set my points as,\[(X_1,Y_1)=(-10,3)\]\[(X_2,Y_2)=(-10,12)\] like i said, as long as you make sure that you follow which set you call as "set 1" and which one you call "set 2" you will get the same answer

OpenStudy (anonymous):

I was doing fine with these equations at first but now Im doing something wrong. Find the distance between the points given: (0,-6) and (9,6) I didn't get any of the choices given for his answer either.

OpenStudy (anonymous):

And for the one with -10, the options are sqrt 17, 9, and 15. @Embryo

OpenStudy (anonymous):

So i'm going to set my points like this\[(X_1,Y_1)=(0,-6)\]\[(X_2,Y_2)=(9,6)\]now when i plug that into the formula, it would look like this\[d=\sqrt{(9-6)^2+(6-(-6))^2}=\sqrt{(3)^2+(6+6)^2}=\sqrt{9+(12)^2}\]so the final solution should be\[\sqrt{9+144}=\sqrt{153}\approx 12.37\]

OpenStudy (anonymous):

yeah, the answer for the one with -10 should be 9

OpenStudy (anonymous):

The one with - 6 has the answers of 9, 15, and sqrt 23 @Embryo

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

oh woops, i accidently used the 6 as my X_1, my bad, it should have looked like this\[\sqrt{(9-0)^2+(6-(-6))^2}=\sqrt{(9)^2+(12)^2}=\sqrt{81+144}=15\]

OpenStudy (anonymous):

I am terribly sorry for that mistake on my part

OpenStudy (anonymous):

It's okay